OpenStudy (anonymous):

Can someone please help me to set this problem up? Thank you. Hamburgers sell for $3.00 each and cheeseburgers sell for $4.00 each. Bob sold a total of 400 burgers at the Marion Popcorn Festival, and he had total sales of $1,450. How many hamburgers did Bob sell? How many cheeseburgers did Bob sell?

OpenStudy (amorfide):

let ham be x let cheese be y 3x+4y=1450 since it is 3dollar per ham 4 dollar per cheese you can get a maximum of 1450 next we know 400 burgers are the maximum you can sell so you have x+y=400 so you get simultaneous equations i believe 3x+4y=1450 x+y=400

OpenStudy (amorfide):

question is, why are they selling burgers a popcorn festival

OpenStudy (anonymous):

I appreciate your help. That is a good question... So much for the general cotton candy & popcorn, etc.

OpenStudy (amorfide):

hope you can solve it from this, if not let me know

OpenStudy (anonymous):

I am going to subtraction 400 for both sides?

OpenStudy (jdoe0001):

well.. maybe the burgers are just corn-burgers and include a free soda and butter, and they're complementing the popcorn =)

OpenStudy (anonymous):

Great idea. ;) Would you be able to help me with this question? I'm unsure how to set it up in order to get both answers of the hamburgers and cheeseburgers.